OpenStudy (anonymous):

In a series circuit with 9V battery and two resistors, one 100 Ohms and one at 350 Ohms, this is the voltage across each one respectively. **What is that voltage?

OpenStudy (michele_laino):

the current in your circuit is: \[\Large I = \frac{{\Delta V}}{{{R_1} + {R_2}}} = \frac{9}{{100 + 350}}\] since your resistors are connected in series, and the equivalent resistance is: R_1+R_2 

OpenStudy (anonymous):

ok! and solving we get 0.02?

OpenStudy (michele_laino):

correct!

OpenStudy (anonymous):

is that the voltage across each one? :/ 0.02 is our solution?

OpenStudy (michele_laino):

now using the Ohm's law, we can compute the requested voltage drop as below:  \[\Large \begin{gathered} \Delta {V_1} = I \times {R_1} = 0.02 \times 100 \hfill \\ \Delta {V_2} = I \times {R_2} = 0.02 \times 350 \hfill \\ \end{gathered} \]